WARNING: This web is not used anymore. Please use PDBService.MinuteS27September07 instead!
 

Minutes 27 September 07

  • Phone:
    • NDGF: Olli
    • PIC: Luis
    • TRIUMF: Denice
    • SARA: Alexander
    • BNL: Carlos
    • ATLAS: Sasha
    • CNAF: Barbara

  • CERN:
    • 3D: Miguel, Dirk, Eva
    • ATLAS: Gancho, Florbela
    • LHCb: Marco
    • COOL: Andrea

  • General announcements:
    • OCP course at CERN: ask interested people at sites to subscribe at web. CERN HR will cancel the course if it does not fill up
    • SAN vs NAS storage:
      • presentation about NAS storage from DES. Very interesting points but better to have a common technology and physics services at T0 use SAN.
      • Question/Answer: no site thinking to move to NAS

  • LFC setup for LHCB setup to be used with streams:
    • CNAF replica have a local copy of ACL tables
    • LHCb prefers to have it centralized at CERN and control the user/group creation and permitions. So ACL tables should be copied to all T1
    • First LFC needs to disable the automatic ACL creation on user first connection. After this the current tables at CERN will be exported to T1 and streams will be keep them synchronized.
    • Needs to be organized will all sites. There will be a wiki with instructions for deployment at sites.
    • Timescale of these changes will be at least 4 weeks.

  • Sites status:
    • CNAF (Barbara):
      • moving LHCb cluster – intervention scheduled for next Wednesday – coordinate with Eva
      • (as the LHCb cluster includes the LFC service, the intervention has been postponed)
    • BNL (Carlos):
      • Floberla has been contacted regarding the resources for the ATLAS experiment.
    • RAL (Gordon):
      • One problem with one of the databases, related to memory, fixed after changing some parameters. Intervention announced
    • BNL and RAL are volunteers for storage for TAGs
    • SARA (Alexander): today, some network problems, looking at them

  • Experiment status:
    • ATLAS:
      • Working plan to see how to get the TAGs on T1 and then on T2
      • Working at CERN first, and then some volunteers to setup the TAGs: BNL, additional: RAL, TRIUMF...
      • Storage for performance tests.
      • Working with ATLAS Tier1 team on the ideas.
      • Florbela will prepare same slides in order to explain the details.
      • COOL: next test: last week of October: first test with PVSS to COOL.
      • Richard tests against CNAF and IN2P3: problem with query: Full table scan on table with CLOBs
      • Andreas has proposed to compare the execution plans using and not using the bind variables (Bind variables for long run queries can be not useful)
      • PVSS streams tests will re-started next week.
      • NDGF is still out. We will try it again next week.
      • We should think on a way to facilitate the COOL data export and imports.

    • LHCb: no more tests have been run.
      • 400 concurrent connections without problems.

  • CERN workshop: Service reliability
    • 26-30 November
    • Organize also the 3d workshop
    • Dates already on the wiki, agenda to be planned
    • This is a proposal, people can send comments
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-10-11 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback